OpenStudy (khalilforthewin):

In a candy box, there are 4 times as many candies with nuts as candies without nuts. There are 25 candies in the box. How many candies with nuts are in the box? 20 5 9 21

OpenStudy (khalilforthewin):

I need help understanding how to get the answer

OpenStudy (anonymous):

25/5=5, 4x5=20

OpenStudy (khalilforthewin):

so its 20?

OpenStudy (anonymous):

\[4x+x=25 \to 5x=25 \to x=5\] So \[4(5) + 5 = 25\] So 20 candies with nuts and 5 without

OpenStudy (khalilforthewin):

how did you get 5

OpenStudy (khalilforthewin):

thank you so much i get it now

OpenStudy (anonymous):

The first section \[4x+x=25 \to 5x=25\] Divide to get x alone so x will be 5. Then just substitute into the original equation.

OpenStudy (anonymous):

And I'm glad you understand :)